# Towards Urban Planing AI Agent in the Age of Agentic AI

### v3

@@ Abstract
Generative AI, large language models, and agentic AI have emerged separately of urban planning. However, the convergence between AI and urban planning presents an interesting opportunity towards AI urban planners. Existing studies conceptualizes urban planning as a generative AI task, where AI synthesizes land-use configurations under geospatial, social, and human-centric constraints and reshape automated urban design. We further identify critical gaps of existing generative urban planning studies: 1) the generative structure has to be predefined with strong assumption: all of adversarial generator-discriminator, forward and inverse diffusion structures, hierarchical zone-POI generative structure are predefined by humans; 2) ignore the power of domain expert developed tools: domain urban planners have developed various tools in the urban planning process guided by urban theory, while existing pure neural networks based generation ignore the power of the tools developed by urban planner practitioners. To address these limitations, we outline a future research direction agentic urban AI planner, calling for a new synthesis of agentic AI and participatory urbanism.
I Introduction and Motivation
Urban planning is an interdisciplinary and complex process that involves with public policy, social science, engineering, architecture, landscape, and other related fields. Effective urban planning can help to mitigate the operational and social vulnerability of a urban system, such as high tax, crimes, traffic congestion and accidents, pollution, depression, and anxiety [1, 2, 3].
@@ I Introduction and Motivation
Recent breakthroughs in Generative AI (GenAI) [4, 5, 6, 7], Large Language Models (LLMs) [8, 9, 10], and Agentic AI [11, 12] offer a transformative opportunity to rethink urban planning. These technologies introduce the abilities to generate alternative futures, understand complex textual and visual data, and act as autonomous agents that reason, simulate, and navigate planning goals. The new abilities of these AI paradigms makes it possible to develop an AI Urban Planner that can augment human expertise, democratize access to planning insights, and adapt to changing urban conditions. This insight prompts a fundamental question: how might AI reshape the practice and ethics of urban planning in the era of AI? What roles does GenAI, LLMs, and agentic AI play in urban planning? Can machines develop and learn at a human capability to automatically and quickly calculate land-use configurations? In this way, machines can be planning assistants and human planners can adjust machine-generated plans for specific needs.
II Existing Studies: Urban Planning as Generative AI
@@
II-A Recent Studies in GenAI for Urban Planning
@@
Drawing on the advanced capabilities of generative AI, we have reenvisioned urban planning as a generative AI task, conceptualizing spatial configuration plans as images, matrices, or tensors, with a generative model tasked with producing the optimal tensor representation [15]. The work in [16] frames urban planning as the generation of land-use configurations over spatially attributed graphs augmented with geographic context. These configurations are encoded as multi-channel tensors representing Point of Interest (POI) distributions based on location and function. A graph autoencoder generates regional embeddings, which are fed into a GAN to assess generated plans using metrics such as willingness to pay, diversity of human activities, and social interaction. Despite its novelty, this approach struggles to model spatial hierarchies and exhibits training instability. In [17], land-use generation is conditioned on spatial context and textual planner guidance through a deep conditional variational autoencoder, yielding zone maps and POI layouts informed by contextual factors like geography and mobility. While effective in integrating planning expertise, it does not explicitly address hierarchical urban structures. To overcome these limitations, [18, 15] introduces a hierarchical conditional transformer that first produces zone-level functions before refining them into detailed grid-level plans. Similarly, [19] formulates planning as a two-tier hierarchical reinforcement learning (RL) problem, employing an actor-critic network for zone-level planning and a Deep Q-Network for POI-level design. Likewise, [20] models cities as graphs, utilizing GNN-based RL to make sequential planning decisions.
II-B The Underlying Generative Urban Planning Framework
@@
Generative urban planning formulates urban planing as conditional generation of land-use configurations, where an AI system learns to synthesize urban layouts that are optimized with respect to spatial, social, and policy-driven objectives. This principled framework consists of two stages: representation learning and conditional generation.
II-B1 Representation: Modeling Urban Contexts and Human Needs
@@
The representation stage aims to encode the multifaceted conditions of a target urban area to plan into a structured embedding that guides the generative model. This involves integrating and representing diverse information: 1) Geospatial Form: Captures the spatial arrangement and morphology of urban environments, including streets, buildings, parks, and infrastructure. These spatial patterns influence accessibility, land value, and neighborhood character. 2) Human Mobility: Derived from data such as GPS traces, public transit logs, and mobile phone signals, mobility data reveals functional flows and commuting behaviors critical to understanding land-use needs. 3) Social Interactions: Encompasses both physical interactions (e.g., in plazas, schools, offices) and digital engagement (e.g., social media activity). These patterns reflect community structure, cultural behavior, and public sentiment. 4) Planner Requirements: Urban planners often express goals—such as increasing green space or housing density—via textual prompts, zoning rules, or categorical preferences. These must be encoded as soft or hard constraints in the model. To unify all these modalities, deep representation learning techniques such as autoencoders, convolutional neural networks (CNNs), and attention-based models are employed to transform high-dimensional inputs into semantically-rich, low-dimensional embeddings. These embeddings serve as a comprehensive representation of a region’s situational context and planning constraints, forming the condition for land-use generation.
II-B2 Generation: Producing Optimal Land-Use and Built Environment Configurations
@@
Given a learned representation, the generative stage synthesizes land-use plans using deep generative models, such as, Variational Autoencoders (VAEs) [21], Generative Adversarial Networks (GANs) [22, 23, 24], Autoregressive Models [25, 26], Flow-Based Models, Energy-Based Models. These generative urban planning models can be trained with a combination of machine learning losses (e.g., reconstruction loss, likelihood maximization, adversarial loss) and application-specific regularization tailored to urban planning. Together, representation and generation form a pipeline where AI systems learn to ”understand” the urban environment and ”create” planning proposals that are not only technically valid but socially and contextually meaningful.
@@
II-C Embedding Generative Urban Planning to Real Systems: Flooding Resilience
@@
One application example is a redevelopment zone along the Gulf Coast, where legacy land-use patterns—such as dense residential clustering near flood-prone areas, fragmented green infrastructure, and spatially isolated critical facilities—have amplified vulnerability to storm surge and inland flooding. The central planning challenge is to reconfigure urban layouts to anticipate, absorb, and recover from flood events, while maintaining economic functionality and social continuity. Generative urban planning offers a promising alternative by integrating AI with spatial design to generate land-use scenarios that are both realistic and adaptive to hydrological risks. Using multimodal data inputs, including flood simulation outputs, infrastructure risk maps, and planner-defined constraints, a conditional generative model can synthesize redevelopment layouts that elevate critical infrastructure, reposition vulnerable housing clusters, introduce green-blue corridors for water absorption, and optimize road networks for emergency access, all while complying with zoning and development guidelines. Such an approach aligns with the emerging ”resilience-by-design” paradigm in urban planning [27].
II-D Limitations of Current AI Urban Planner Research
@@
Despite advances in AI for urban planning, current approaches face critical limitations across five foundational dimensions: theory, methodology, data, computing, and applications. 1) Theory-wise, most AI models reduce planning to spatial optimization or land-use generation, thus, ignore core urban concepts such as zoning, resilience, social equity, and participatory governance, and often ignore the political and normative nature of planning decisions—undermining interpretability and legitimacy. 2) Method-wise, models [16, 20, 19, 18, 15, 17, 28] rely on single-step or loosely staged generation conditioned on static inputs that limit creativity and adaptivity, fail to capture multi-granularity dynamics, multiple planning objective trade-offs, and adaptive processes. 3) Data-wise, existing city-specific datasets (e.g., Beijing [29], NYC) lacks clear evaluation criteria on scoring a good planning and a bad planning. 3) Computation-wise, GenAI models are resource-intensive and prone to instability, especially in adversarial or variational settings; they lacks explainability and transparency due to the use of neural network based architectures, and often remain opaque to end users. 4) Application-wise, existing generative planning methods ignore many domain expert developed tools. Be sure to notice that domain urban planners have developed various tools to assist each step of urban planning process guided by planning theory, while existing pure neural networks based generation ignore the power of the tools developed by urban planner practitioners.
@@
III The Vision: Integrating VLM, Multimodal Reasoning, and Agentic AI for Urban Planning AI Agents
@@
To go beyond GenAI like VAEs, GANs, or diffusion models, we believe that VLMs integrate visual and textual modalities and enable more semantically grounded, instruction-following generation [30, 31]. For example, we can align urban design prompts (e.g., “generate a transit-oriented mixed-use layout”) with satellite imagery, GIS data, and street-level visuals, in order to allow planners to interact with models using natural language while preserving spatial fidelity. Agentic AI further models the complex urban planning pipeline by introducing planning, reasoning, task decomposition, and multi-step decision-making abilities. Agent-based urban planners can autonomously explore spatial trade-offs, simulate long-term effects, and iteratively refine plans based on feedback, environmental constraints, or human instructions. VLMs and agentic systems enable an AI urban planner that is not only generative, but goal-driven, context-aware, and aligned with multimodal planning tasks that are capable of understanding complex instructions, grounding outputs in real-world data, and dynamically adapting to evolving objectives.
@@
Despite the interdisciplinarity and complex nature of urban planning, arising from its multifaceted requirements, agentic AI introduces a new paradigm that directly responds to these challenges. The integration of reasoning, task decomposition, and domain-specific tool usage empowers agentic AI to revolutionize urban planning towards more transparent, explainable, and practical urban configurations.
Reasoning plays a crucial role in this paradigm, enabling agentic AI to go beyond static generation.By systematically analyzing the diverse mixture of planning objectives, constraints, and potential conflicts, an AI urban planner can apply logical and context-aware reasoning to explicitly balance trade-offs across multiple urban planning requirements, thereby facilitating informed decision-making, anticipating potential downstream consequences, and producing planning outcomes that are both technically feasible and contextually meaningful.
Task decomposition enables agentic AI to effectively address the inherent complexity of urban planning. Instead of generating a comprehensive urban planning alternative in a single step, agentic AI system adopts a divide-and-conquer paradigm that decomposes the planning process into manageable sub-tasks, where each sub-task focuses on a specific dimension of the planning. Importantly, each sub-task can also be governed by its own objectives and constraints, allowing the system to incorporate local requirements while maintaining consistency at the system level, ensuring more transparent and traceable intermediate decisions.
@@
Domain-specific tool usage further enhances agentic AI by integrating practical tools developed by urban planning practitioners. These tools have been repeatedly validated in real-world projects and regulatory processes, including decades of planning theory and professional practice. By leveraging such domain-expert designed tools, agentic AI system can align the outputs with up-to-date institutional standards and long-standing urban planning principles rather than relying solely on static data-driven inference, facilitating innovative, credible and practically applicable urban planning alternatives.These capabilities form a comprehensive pipeline that effectively analyzes, summarizes, and generates optimal planning decisions under evolving urban objectives, thereby addressing the inherent complexity of urban planning and rendering agentic AI systems more practical and advanced.

### v4

# Urban Planning in the Age of Agentic AI: Emerging Paradigms and Prospects

@@ Abstract
Generative AI, large language models (LLMs), and agentic AI have emerged separately of urban planning. However, the convergence between AI and urban planning presents an interesting opportunity towards AI urban planners. Existing studies conceptualizes urban planning as a generative AI task, where AI synthesizes land-use configurations under geospatial, social, and human-centric constraints and reshape automated urban design. We further identify critical gaps of existing generative urban planning studies: 1) the generative structure has to be predefined with strong assumption: all of adversarial networks, diffusion models, hierarchical zone-POI generative structure are predefined by humans; 2) ignore the power of domain expert developed tools: domain urban planners have developed various tools in the urban planning process guided by urban theory, while existing pure neural networks based generation ignore the power of the tools developed by urban planner practitioners. To address these limitations, we outline a future research direction agentic urban AI planner, calling for a new synthesis of agentic AI and participatory urbanism that integrates AI capabilities with domain expertise and public engagement.
1 Introduction and Motivation
Urban planning is an interdisciplinary and complex process that involves public policy, social science, engineering, architecture, landscape, and other related fields. Effective urban planning can help to mitigate the operational and social vulnerability of a urban system, such as high tax, crimes, traffic congestion and accidents, pollution, depression, and anxiety [1, 2, 3].
@@ 1 Introduction and Motivation
1Recent breakthroughs in Generative AI (GenAI) [4, 5, 6, 7], Large Language Models (LLMs) [8, 9, 10], and Agentic AI [11, 12] offer a transformative opportunity to rethink urban planning. These technologies introduce the abilities to generate alternative futures, understand complex textual and visual data, and act as autonomous agents that reason, simulate, and navigate planning goals. The new abilities of these AI paradigms make it possible to develop an AI Urban Planner that can augment human expertise, democratize access to planning insights, and adapt to changing urban conditions. This insight prompts a fundamental question: how might AI reshape the practice and ethics of urban planning in the era of AI? What roles does GenAI, LLMs, and agentic AI play in urban planning? Can machines develop and learn at a human capability to automatically and quickly calculate land-use configurations? In this way, machines can be planning assistants, while human planners can adjust machine-generated plans for specific needs.
2 Existing Studies: Urban Planning as Generative AI
@@
2.1 Recent Studies in GenAI for Urban Planning
@@
Drawing on the advanced capabilities of generative AI, we have reenvisioned urban planning as a generative AI task, conceptualizing spatial configuration plans as images, matrices, or tensors, with a generative model tasked with producing the optimal tensor representation [15] under specific geographic and social constraints. The work in [16] frames urban planning as the generation of land-use configurations over graphs enriched with geographic attributes and contextual information. These configurations are encoded as multi-layered data representations representing Point of Interest (POI) distributions based on location and function. A graph autoencoder generates regional embeddings, which are fed into a GAN to assess generated plans using metrics such as willingness to pay, diversity of human activities, and social interaction. Despite its novelty, this approach struggles to model spatial hierarchies and exhibits training instability. In [17], land-use generation is conditioned on spatial context and textual planner guidance through a deep conditional variational autoencoder, yielding zone maps and POI layouts informed by contextual factors like geography and mobility. While effective in integrating planning expertise, it does not explicitly address hierarchical urban structures. To overcome these limitations, [18, 15] introduces a hierarchical conditional transformer that first produces zone-level functions before refining them into detailed grid-level plans. Similarly, [19] formulates planning as a two-tier hierarchical reinforcement learning (RL) problem, employing an actor-critic network and a Deep Q-Network to support decision-making at the zone and POI levels, respectively. Likewise, [20] models cities as graphs, utilizing GNN-based RL to make sequential planning decisions.
2.2 The Underlying Generative Urban Planning Framework
@@
Generative urban planning formulates urban planing as conditional generation of land-use configurations, where an AI system learns to synthesize urban layouts that are optimized with respect to spatial, social, and policy-driven objectives. This principled framework consists of two stages: learning representations of urban contexts and needs and generating land-use scenarios under those conditions.
2.2.1 Representation: Modeling Urban Contexts and Human Needs
@@
The representation stage aims to encode the multifaceted conditions of a target urban area to plan into a structured embedding that guides the generative model. This involves integrating and representing diverse information: 1) Geospatial Form: Captures the spatial arrangement and morphology of urban environments, including streets, buildings, parks, and infrastructure. These spatial patterns influence accessibility, land value, and neighborhood character. 2) Human Mobility: Derived from data such as GPS traces, public transit logs, and mobile phone signals, mobility data reveals functional flows and commuting behaviors critical to understanding land-use needs. 3) Social Interactions: Encompasses both physical interactions (e.g., in plazas, schools, offices) and digital engagement (e.g., social media activity). These patterns reflect community structure, cultural behavior, and public sentiment. 4) Planner Requirements: Urban planners often express goals—such as increasing green space or housing density—via textual prompts, zoning rules, or categorical preferences. These must be encoded as flexible or strict constraints in the model. To unify all these modalities, deep representation learning techniques such as autoencoders, convolutional neural networks (CNNs), and attention-based models are employed to transform high-dimensional inputs into semantically-rich, low-dimensional embeddings. These embeddings serve as a comprehensive representation of a region’s situational context and planning constraints, forming the condition for land-use generation.
2.2.2 Generation: Producing Optimal Land-Use and Built Environment Configurations
@@
Given a learned representation, the generative stage synthesizes land-use plans using deep generative models, such as, Variational Autoencoders (VAEs) [21], Generative Adversarial Networks (GANs) [22, 23, 24], Autoregressive Models [25, 26], and other generative techniques. These generative urban planning models can be trained with a combination of machine learning objectives and planning-specific constraints, ensuring that generated designs remain both technically robust and aligned with urban planning principles. Together, representation and generation form a pipeline where AI systems learn to ”understand” the urban environment and ”create” planning proposals that are not only technically valid but socially and contextually meaningful.
@@
2.3 Embedding Generative Urban Planning to Real Systems: Flooding Resilience
@@
One application example is a redevelopment zone along the Gulf Coast, where legacy land-use patterns—such as dense residential clustering near flood-prone areas, fragmented green infrastructure, and spatially isolated critical facilities—have amplified vulnerability to storm surge and inland flooding. The central planning challenge is to reconfigure urban layouts to anticipate, absorb, and recover from flood events, while maintaining economic functionality and social continuity. Generative urban planning offers a promising alternative by integrating AI with spatial design to generate land-use scenarios that are both realistic and adaptive to hydrological risks, offering planners new options beyond conventional rule-based methods. Using multimodal data inputs, including flood simulation outputs, infrastructure risk maps, and planner-defined constraints, a conditional generative model can synthesize redevelopment layouts that elevate critical infrastructure, reposition vulnerable housing clusters, introduce green-blue corridors for water absorption, and optimize road networks for emergency access, all while complying with zoning and development guidelines. Such an approach aligns with the emerging ”resilience-by-design” paradigm in urban planning [27].
2.4 Limitations of Current AI Urban Planner Research
@@
Despite advances in AI for urban planning, current approaches face critical limitations across five foundational dimensions: theory, methodology, data, computing, and applications. 1) Theory-wise, most AI models reduce planning to spatial optimization or land-use generation, thus, ignore core urban concepts such as zoning, resilience, social equity, and participatory governance, and often ignore the political and normative nature of planning decisions—undermining interpretability and legitimacy. 2) Method-wise, models [16, 20, 19, 18, 15, 17, 28] rely on single-step or loosely staged generation conditioned on static inputs that limit creativity and adaptivity, fail to capture multi-granularity dynamics, multiple planning objective trade-offs, and adaptive processes. 3) Data-wise, existing city-specific datasets (e.g., Beijing [29], NYC) lack clear evaluation criteria on scoring a good planning and a bad planning. 4) Computation-wise, GenAI models are resource-intensive and prone to instability, especially in adversarial or variational settings; they lacks explainability and transparency due to the use of neural network based architectures, and often remain opaque to end users. 5) Application-wise, existing generative planning methods ignore many domain expert developed tools. Be sure to notice that domain urban planners have developed various tools based on planning theory to assist each step of urban planning process, while existing pure neural networks based generation overlooks the power of these established practices.
@@
3 The Vision: Integrating VLM, Multimodal Reasoning, and Agentic AI for Urban Planning AI Agents
@@
To go beyond GenAI like VAEs, GANs, or diffusion models, we believe that Vision-Language Models (VLMs) integrate visual and textual modalities to produce outputs that are semantically meaningful and instruction-following [30, 31]. This multimodal capacity allows AI systems to integrate maps, imagery, and policy goals to pursue complex planning objectives. For example, urban design prompts (e.g., “generate a transit-oriented mixed-use layout”) can be aligned with satellite imagery, GIS data, and street-level visuals. This allows planners to interact with models using natural language, while ensuring that outputs remain spatially accurate and grounded in real-world data. Agentic AI extends this capability by modeling the planning pipeline itself, with abilities for reasoning, task decomposition, and multi-step decision-making that mirror how human planners approach complex planning tasks. Agent-based urban planners can autonomously explore spatial trade-offs, simulate long-term effects, and iteratively refine plans based on feedback, environmental constraints, or guidance from human planners. Together, VLMs and agentic systems enable an AI urban planner that is not only generative but also goal-driven and context-aware. Such AI systems can understand complex instructions, align outputs with multimodal data, and dynamically adapt proposals to evolving objectives.
3.1 Agentic AI Abilities: Implications for Urban Planning
@@
Agentic AI introduces a new paradigm for urban planning by directly addressing its interdisciplinary and complex nature. Through the integration of reasoning, task decomposition, and the use of domain-specific tools, agentic AI has the potential to make planning processes more transparent, explainable, and practical.
3.1.1 Reasoning
Unlike static models that simply generate outputs, agentic AI can systematically analyze diverse objectives, constraints, and potential conflicts. This reasoning ability enables the AI urban planners to balance objectives such as resilience, equity, and cost, while highlighting trade-offs and anticipating downstream consequences—making planning decisions more transparent and better aligned with real-world needs.
3.1.2 Task decomposition
Urban planning is too complex to be solved in a single step. Agentic AI addresses this challenge with a divide-and-conquer approach, decomposing the process into smaller, manageable tasks, each focused on a particular planning dimension such as land use, mobility, or infrastructure. These sub-tasks can follow their own objectives and constraints while remaining consistent with the overall plan, making the decision process more transparent, traceable, and easier for planners to engage with.
3.1.3 Domain-specific tool usage
@@
Another strength of agentic AI is its ability to integrate practical tools developed by urban planning practitioners. These tools (e.g., simulation software, zoning analysis methods, resilience metrics) have been repeatedly validated in real-world projects and regulatory processes, and are grounded in decades of planning theory and professional practice. By leveraging such domain-expert tools, agentic AI can align its outputs with institutional standards and established planning principles rather than relying solely on static, data-driven inference. This makes the results more innovative, credible, and directly applicable to practice. Through the integration of these abilities, agentic AI emerges as a practical paradigm for urban planning. It can handle complex and conflicting objectives in a transparent and traceable way, while aligning outputs with established planning standards and producing results that are credible to experts and accessible to stakeholders.